OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (raffle_snaffle):

what is the formula for the equation of a line?

OpenStudy (anonymous):

im not sure:/ @raffle_snaffle

OpenStudy (raffle_snaffle):

y=mx+b

OpenStudy (anonymous):

so I would multiply the slope by x+b?

OpenStudy (anonymous):

what does b stand for?

OpenStudy (anonymous):

@raffle_snaffle

OpenStudy (raffle_snaffle):

do you remember the point slope formula?

OpenStudy (anonymous):

y - y1 = m(x - x1)

OpenStudy (raffle_snaffle):

use that forumula to find the equation of a line=)

OpenStudy (anonymous):

okay, thank you

OpenStudy (raffle_snaffle):

yeah np

OpenStudy (raffle_snaffle):

use that forumula to find the equation of a line=)

OpenStudy (raffle_snaffle):

do you remember the point slope formula?